# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Medical Image

### v1

@@ Abstract
Semantic medical image segmentation is a crucial part of both scientific research and clinical care. With enough labelled data, deep learning models can be trained to accurately automate specific medical image segmentation tasks. However, manually segmenting images to create training data is highly labor intensive. In this paper, we present ScribblePrompt, an interactive segmentation framework for medical imaging that enables human annotators to segment unseen structures using scribbles, clicks, and bounding boxes. Scribbles are an intuitive and effective form of user interaction for complex tasks, however most existing methods focus on click-based interactions. We introduce algorithms for simulating realistic scribbles that enable training models that are amenable to multiple types of interaction. To achieve generalization to new tasks, we train on a diverse collection of 65 open-access biomedical datasets – using both real and synthetic labels. We test ScribblePrompt on multiple network architectures and unseen datasets, and demonstrate that it can be used in real-time on a single CPU. We evaluate ScribblePrompt using manually-collected scribbles, simulated interactions, and a user study. ScribblePrompt outperforms existing methods in all our evaluations. In the user study, ScribblePrompt reduced annotation time by 28% while improving Dice by 15% compared to existing methods. We showcase ScribblePrompt in an online demo and provide code at scribbleprompt.csail.mit.edu.
Figure 1: Example predictions with iterative simulated interactions (steps). ScribblePrompt enables iterative interactive segmentation of unseen tasks using scribbles, clicks, and bounding boxes. We show representative examples with simulated interactions that reflect the diversity of tasks in our evaluation data. At each step, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue.
@@ 1 Introduction
Semantic segmentation of medical images is an essential step in a wide range of biomedical research and clinical care pipelines. Deep learning has become the predominant method to automate existing segmentation tasks [108, 53, 123, 94]. However, training accurate models is a large undertaking that depends on having diverse images carefully annotated by skilled experts. Biomedical researchers and clinicians often target novel segmentation tasks, requiring the delineation of either new regions of interest [142] or new imaging modalities [113, 91].
Most medical image segmentation systems used in practice provide simplistic editing tools, sometimes with intensity-based algorithmic assistance, such as ITK-SNAP [141]. Recent research has focused on learning-based interactive segmentation systems in which the annotator iteratively collaborates with a model to complete the segmentation, providing prompts, such as clicks, and receiving predicted segmentation(s) that they can correct with additional prompting.
Given the diversity of biomedical imaging available, interactive segmentation tools would be particularly useful when able to segment previously unsegmented biomedical domains, regions of interest, and image types. However, existing methods in this space have not been widely adopted in practice, mostly because they are useful only in limited circumstances or for specific segmentation targets such as abdominal organs in CT [80, 111, 131, 100, 130, 150]. Recent foundation models tackle more general purpose semantic segmentation [61, 151, 132, 133], but also do not generalize to new medical imaging tasks without fine-tuning for the desired tasks [90, 50, 44, 117, 82, 25, 135]. Other methods avoid fine-tuning but require extensive “prompts” involving already labeled examples [20, 28].
Most prior work on interactive segmentation has been limited to click-based [120, 76, 23, 13, 80, 61, 137, 70, 34] and bounding box-based interactions [138, 61, 106], often requiring users to follow specialized click protocols [144] such as clicking extreme points [85] or interior margin points [80]. Scribbles are a natural mode of interaction; a large, rough, scribble is often easier and faster to perform than a series of carefully placed clicks. Scribbles provide more information than clicks, which is helpful for segmenting complex objects, with minimal increase in user effort. Unfortunately, few existing methods employ scribbles, because of the challenges of acquiring or simulating realistic scribbles for training and evaluation [24].
@@
In this paper, we present ScribblePrompt, an interactive segmentation framework that:
  • enables scribble inputs (along with clicks and bounding boxes), to give users more flexibility. To train the model, we devise ways to simulate diverse and realistic scribbles.
  • enables generalization to unseen segmentation tasks and medical imaging datasets. In addition to training on a large collection of datasets, we propose augmentation strategies using synthetic labels.
  • facilitates interactive refinement of predictions. We simulate iterative interactions during training, and train a single network to handle both the initial prediction and refinement of predictions based on user corrections.
@@
  • is highly efficient, enabling real-time inference, even on a single CPU.
We demonstrate that ScribblePrompt is more accurate than existing methods when tested on manually-collected scribbles and simulated interactions (Figure 1) on twelve datasets. Through a user study with experienced annotators, we show that ScribblePrompt’s ability to segment accurately based on bounding boxes, clicks and scribbles provides valuable flexibility to users, reducing annotation time by 28% while increasing Dice by 15% compared to previous methods.
@@
We release an interactive tool, code, and model weights at scribbleprompt.csail.mit.edu. We will also release new multi-annotator segmentation labels created using ScribblePrompt for select datasets.
@@ 2 Related Works
Interactive Medical Image Segmentation. Deep learning-based techniques for interactive segmentation of medical images [80, 100, 130, 131] have gained popularity in research over traditional intensity-based methods [141, 40, 17, 27, 126]. Many existing methods focus on using human interaction to improve the accuracy of segmentation for tasks where some labelled data exists, but fully-automatic segmentation methods fail to produce accurate segmentations [150, 131, 129, 80]. These methods assume that the model will be used to perform the same segmentation task(s) it was trained on, leading to domain-specific solutions [131, 7, 75, 96, 106, 129]. A few methods generalize to new, but similar, classes [111] and modalities [80, 130] to those seen in training. In contrast, we design ScribblePrompt to help annotators segment a wide range of medical images at inference time.
@@
Generalist Models. Recent vision foundation models employ prompting to reduce the burden of manual segmentation and enable generalization to new tasks. These models, trained on large collections of natural image data, segment potentially-unseen structures specified by spatial prompts [61, 151], text descriptions [104, 61, 151], or examples [132, 133, 20, 28]. Some of these models [61, 151] can be used for interactive segmentation, however they are optimized for performance on natural imaging tasks.
@@
Fine-tuning models initially developed using natural images is often unhelpful for specific tasks in the medical imaging domain [105]. The Segment Anything Model [61], a foundation model trained on natural images, performs well at medical image segmentation tasks with clear boundaries (e.g., organ segmentation in abdominal CT), but performs poorly on tasks and modalities involving more subtle boundaries (e.g., deep structures in Brain MRI) [50, 90, 117, 44]. Several recent or concurrent papers fine-tune SAM for specific tasks or modalities [97, 73, 128, 58, 49, 143]. A few fine-tune SAM to segment medical images from multiple-modalities with bounding boxes [82], clicks [135] or both [25]. However, as we show in our experiments, these models don’t perform well on many unseen tasks, including those with thin structure (e.g., blood vessels).
@@
Several natural imaging methods simulate iterative user interactions and condition the model on the previous predictions to train a single network to make the initial prediction and perform refinement [120, 61, 119, 18, 34, 76]. This approach is appropriate for natural images where there is ample data and typically one modality. However, such methods have not previously been developed for the medical imaging domain, which is more fragmented because of specialized regions of interest and different modalities. We assemble a large diverse collection of medical imaging datasets and develop models that facilitate iterative interactive segmentation of unseen medical imaging tasks.
@@
User Interaction. Users can prompt interactive segmentation systems in many different ways, such as bounding boxes [110, 139, 61, 13, 130, 106], scribbles [131, 24, 129], and clicks [131, 80, 13, 76, 120]. Few interactive segmentation models incorporate multiple types of inputs. Some works have explored training with more specialized types of clicks [144], such as extreme points [85, 109], interior margin points [80] and center clicks [120, 137] to maximize the information per interaction. However, specialized interactions often require more user time, and models trained only with such interactions are less robust to deviations from the interaction protocol [139, 120, 13]. In contrast, we develop a user-friendly model that performs well under a variety of interaction scenarios.
@@
Scribbles are an intuitive form of interaction, however few interactive segmentation works utilize them, because simulating realistic scribbles is challenging. Many intensity-based image segmentation methods that do not require training can predict segmentations from scribbles, however they require extensive interactions to produce accurate segmentations [40, 17]. Previous deep-learning models trained for scribble-based segmentation either relied on collecting large datasets of manually drawn scribbles [150, 71] or simplistic approaches such as sampling random points [9, 6, 131] or Bezier curves [3]. One work explored simulating more complex scribbles for natural image segmentation [24]. We build on these concepts to create a new scribble simulation engine for training and evaluating scribble-based segmentation methods.
3 ScribblePrompt Method
@@
We present an interactive segmentation method that can generalize to new medical modalities and regions of interest. We describe the problem formulation, and present the important aspects of the ScribblePrompt framework: (i) simulation of interactions during training, (ii) augmentation with synthetic labels during training to encourage generalization, and (iii) an efficient architecture for real-time interaction.
@@ 3.1 Problem Formulation
Let $t$ be a segmentation task consisting of an image and segmentation pair, $\{x^{t},y^{t}\}$. At step $i$, given an image $x^{t}$, a set of user interactions $u_{i}$, and previous prediction $\hat{y}^{t}_{i-1}$, we learn function $f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$ that produces a segmentation $\hat{y}_{i}$ and has learnable parameters $\theta$. The set of interactions $u_{i}$, which may include positive or negative scribbles, positive or negative clicks, and bounding boxes, is provided by a user who has access to the previous prediction, $\hat{y}_{i-1}^{t}$.
We minimize the difference between $y^{t}$ and the series of iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$,
$\small\mathcal{L}(\theta;\tau)=\mathbb{E}_{t\in\tau}\left[\mathbb{E}_{(x^{t},y^{t})\in t}\left[\sum_{i=1}^{k}\mathcal{L}_{Seg}\left(y^{t},f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})\right)\right]\right]\,$ (1)
@@
During training, we sample a task $t$ from training data collection $\tau$, an image $x^{t}$, and segmentation map $y^{t}$. We simulate a possible set of interactions $u_{i}$ based on $y^{t}$, and predict $\hat{y}_{i}^{t}=f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$. We simulate the next set of interactions $u_{i+1}$ based on the error region $\varepsilon_{i}^{t}=y^{t}-\hat{y}_{i}^{t}$, and repeat for $k$ iterations. In the following sections, we describe strategies for simulating $u_{i}$, collecting $\tau$, sampling $(x^{t},y^{t})$, building $f_{\theta}$ and optimizing $\mathcal{L}(\theta;\tau)$.
Figure 2: Simulated scribbles and clicks. Positive interactions (green) are simulated on the segmentation label $y^{t}$ (blue), while negative interactions (red) are simulated on the background $1-y^{t}$.
@@ 3.2 Prompt Simulation
To develop robustness to user interactions $u_{i}$, we introduce algorithms for simulating scribbles, clicks, and bounding box inputs during training. Each scribble and click strategy (Figure 2) can be applied to a ground truth segmentation label $y^{t}$ to simulate positive interactions, or to the background $1-y^{t}$ to simulate negative interactions. We simulate positive and negative correction scribbles or clicks by applying the same strategies to the false negative region $\varepsilon_{i}^{t}>0$ and false positive region $\varepsilon_{i}^{t}<0$ for error region $\varepsilon_{i}^{t}$.
Scribbles. Given binary mask $y\in\{0,1\}^{h\times w}$, we simulate a scribble mask $s\in[0,1]^{h\times w}$ by first generating clean scribbles, and then corrupting them. We illustrate this process in Appendix B.1.1. We start with either:
@@
    Line Scribbles: We draw random lines by sampling two end points from $\{(u,v)|y_{uv}=1\}$.
@@
    Centerline Scribbles: We simulate scribbles drawn in the center of the label. We use a thinning algorithm [145] to draw a 1-pixel wide skeleton of $y$.
@@
    Contour Scribbles: We simulate a rough contour of the desired segmentation within the boundaries of the mask. We first blur the reverse mask $\overline{y}=1-y$ to reduce the size of the label such that $\tilde{y}:=1-\max(\overline{y},\overline{y}\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{y}<h$ sampled in some intensity range $h\sim U[\tilde{y}_{min},\tilde{y}_{max}]$ and use an edge detection algorithm to extract a contour inside the boundary of the mask.
To limit the size of the centerline and contour scribbles, we apply a random mask that breaks the scribbles into smaller parts. We generate the random mask by sampling a smooth noise image $p$, where each pixel is sampled independently from $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$, applying Gaussian blur, and then thresholding it at $\mu_{p}$.
Then, we warp the scribble mask $s$ using a random deformation field $\phi$ to vary the scribble shape and thickness. We ensure the resulting scribble is consistent with mask $y$, by multiplying the warped scribble mask $s\circ\phi$ by $y$.
Clicks. Given a binary mask $y\in\{0,1\}^{h\times w}$, we simulate $n\sim U[n_{min},n_{max}]$ clicks at a time using one of three strategies illustrated in Figure 2:
@@
    Center clicks: We place clicks at the center of each component of the label. First, we create a multi-label mask $m\in\{1,\dots,C\}^{h\times w}$ identifying the $C$ disconnected components of label mask $y$. We then sample pixel coordinates from the set of center points of the components based on the euclidean distance transform, similar to [120, 137].
@@
    Interior border region clicks: We sample clicks from a border region inside the boundary of the mask. We first blur the reverse mask $\overline{y}=1-y$ to reduce the size of the label such that $\tilde{y}:=1-\max(\overline{y},\overline{y}\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{y}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{y}_{min},\tilde{y}_{max})$ are thresholds sampled in some intensity range. We illustrate this process in Appendix B.1.2.
Since users are inclined to spread out their clicks, we impose a minimum separation of a few pixels between random clicks and border region clicks sampled in the same step.
Bounding Boxes. We compute the minimum bounding box that encloses the label $y^{t}$, and enlarge each dimension by $r\sim U[0,20]$ pixels to account for human error.
Iterative Training. During the first step ($i=1$) we sample the combination of interactions and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The Initial set of interactions $u_{1}$ are simulated based on the ground truth label $y^{t}$. In subsequent steps, correction scribbles or clicks are sampled from the error region between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth segmentation $y^{t}$. Since a user can make multiple corrections before refreshing the prediction, during training we sample $n_{cor}\sim U[n_{min},n_{max}]$ correction prompts (scribbles or clicks) per step.

### v2

# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Biomedical Image

@@ Abstract
Biomedical image segmentation is a crucial part of both scientific research and clinical care. With enough labelled data, deep learning models can be trained to accurately automate specific biomedical image segmentation tasks. However, manually segmenting images to create training data is highly labor intensive and requires domain expertise. We present ScribblePrompt, a flexible neural network based interactive segmentation tool for biomedical imaging that enables human annotators to segment previously unseen structures using scribbles, clicks, and bounding boxes. Through rigorous quantitative experiments, we demonstrate that given comparable amounts of interaction, ScribblePrompt produces more accurate segmentations than previous methods on datasets unseen during training. In a user study with domain experts, ScribblePrompt reduced annotation time by 28% while improving Dice by 15% compared to the next best method. ScribblePrompt’s success rests on a set of careful design decisions. These include a training strategy that incorporates both a highly diverse set of images and tasks, novel algorithms for simulated user interactions and labels, and a network that enables fast inference. We showcase ScribblePrompt in an interactive demo and provide code athttps://scribbleprompt.csail.mit.edu
Figure 1: ScribblePrompt enables rapid iterative interactive segmentation of unseen tasks using bounding boxes, clicks, and scribbles. We show predictions from ScribblePrompt with iterative interaction steps on examples from datasets unseen during training. At each step, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue. Scribble thickness is enlarged for visual clarity. See Appendix for more examples.
@@ 1 Introduction
Biomedical image segmentation is an essential step in a wide range of biomedical research and clinical care pipelines. Deep learning has become the predominant method to automate existing segmentation tasks [111, 55, 126, 96]. Biomedical researchers and clinicians often encounter novel segmentation tasks involving either new regions of interest [144] or new image modalities [116, 93]. Unfortunately, supervised training of accurate models for new domains requires diverse images with careful annotations by skilled experts.
Most widely used interactive segmentation systems for biomedical imaging provide minimal, intensity-based, algorithmic assistance [143]. Despite a growing literature, learning-based systems are less widely used in practice, perhaps because they are generally focused on specific tasks or modalities, which limits their applicability [82, 114, 134, 102, 133, 152]. Recent vision foundation models target broad applicability, but require fine-tuning to the medical domain [63, 153, 135, 136]. Most interactive segmentation models require specific interactions, like carefully placed clicks, making them easy to develop but difficult to use in practice. The result of these shortcomings is that existing learning-based models are not widely used in practice.
@@
In this work, we present ScribblePrompt, a general model for interactive biomedical image segmentation that
  1. 1.
    enables users to rapidly and accurately accomplish any biomedical image segmentation task, outperforming state-of-the-art models, particularly for unseen labels and image types;
  2. 2.
    is flexible to different annotation styles, including bounding boxes, clicks, and scribbles,
  3. 3.
@@
    is computationally efficient, enabling real-time inference, even on a single CPU.
To achieve this, we focused on design decisions aimed at ease of use and realistic interactions. To train our model, we started by gathering a large corpus of biomedical imaging datasets. We then designed task augmentation through synthesis strategies to encourage generalization to unseen tasks. Finally, we designed new interaction simulation strategies for training. The model itself uses an architecture optimized for efficient inference, to ensure real-time utility.
To evaluate ScribblePrompt, we compared its performance to that of existing interactive segmentation methods on datasets unseen during training. These experiments involved manually-collected scribbles, simulated interactions (Fig. 1), and a user study in which experienced annotators were asked to segment images. In the user study, ScribblePrompt reduced annotation time by 28% while increasing Dice by 15% compared to the next best method, the Segment Anything Model (SAM) [63]. Given similar amounts of interaction, ScribblePrompt achieved consistently better Dice scores than other methods.
@@
We release an interactive tool, code, model weights, and our dataset of manually-collected scribbles at scribbleprompt.csail.mit.edu. We will also release new multi-annotator segmentation labels created using ScribblePrompt for select datasets.
@@ 2 Related Works
Interactive Biomedical Image Segmentation. Early research into interactive segmentation of biomedical images focused on traditional intensity-based methods [143, 41, 17, 26, 129]. Recent deep-learning based techniques  [82, 102, 133, 134] use human interaction to improve the accuracy of segmentation tasks where labelled data exists, but fully-automatic segmentation methods fail to produce accurate segmentations. This approach assumes that the model will be used to perform the same segmentation task(s) it was trained on, leading to domain-specific solutions [134, 7, 77, 98, 108, 132]. A few methods generalize to new, but similar, classes [114] and modalities [82, 133] to those seen in training. In contrast, ScribblePrompt is designed to help annotators segment a wide range of biomedical images and (potentially unseen) tasks at inference time.
@@
Foundation Models. Recent vision foundation models employ prompting to enable generalization to new tasks. These models, trained on large collections of (natural) image data, segment potentially-unseen structures specified by spatial prompts [63, 153], text descriptions [106, 63, 153], or examples [135, 136, 20, 27]. Some of these models can be used for interactive segmentation with spatial prompts.
@@
Fine-tuning models initially developed using natural images is often unhelpful for specific tasks in the biomedical imaging domain, compared to training from scratch [107]. The Segment Anything Model (SAM) [63], a foundation model trained on natural images, performs well at biomedical image segmentation tasks with clear boundaries (e.g. organ segmentation in abdominal CT), but performs poorly on tasks and modalities involving more subtle delineations (e.g. deep structures in Brain MRI) [51, 92, 120, 45]. Several recent or concurrent papers fine-tune SAM for specific biomedical tasks or modalities [99, 75, 131, 60, 50, 145]. A few fine-tune SAM to segment biomedical images from multiple-modalities with bounding boxes [84], clicks [138] or both [24]. We show in our experiments that these models don’t perform well on many unseen tasks and are limiting in their required interactions.
@@
Several natural image methods simulate iterative user interactions and condition the model on previous predictions to train a single network to make the initial prediction and perform refinement [123, 63, 122, 18, 35, 78]. This approach is appropriate for natural images where there is ample data and typically one modality. However, such methods have not previously been developed for the biomedical imaging domain, which is more fragmented because of specialized regions of interest and different modalities. We use iterative ideas in ScribblePrompt, but we facilitate iterative interactive segmentation of unseen biomedical imaging tasks.
@@
User Interaction. Users can prompt interactive segmentation systems in many different ways, such as bounding boxes [113, 141, 63, 13, 133, 108], scribbles [134, 23, 132], or clicks [134, 82, 13, 78, 123]. Few interactive segmentation models incorporate multiple types of inputs. Some works have explored training with more specialized types of clicks [146], such as extreme points [87, 112], interior margin points [82], and center clicks [123, 140], to maximize the information per interaction. Specialized interactions often require more user time, and models trained only with such interactions are less robust to deviations from the interaction protocol [141, 123, 13]. In contrast, we focus on a user-friendly model that performs well under a variety of interaction scenarios.
@@
Scribbles are an intuitive form of interaction, however few interactive segmentation works employ them, because acquiring realistic scribbles for training and evaluation is challenging. Previous deep-learning models trained for scribble-based segmentation either relied on collecting large datasets of manually drawn scribbles [152, 73, 70] or simplistic approaches such as sampling random points [9, 6, 134] or Bezier curves [3]. One work has explored simulating more complex scribbles for natural image segmentation [23]. We build on these concepts to create a new scribble simulation engine for training and evaluation.
3 ScribblePrompt Approach
@@
We present an interactive segmentation method that can generalize to new biomedical imaging modalities and regions of interest, while staying focused on practical usability. We describe the problem formulation, and present the important aspects of the ScribblePrompt framework: (i) simulation of realistic and flexible interactions during training, (ii) augmentation with synthetic labels during training to encourage generalization, and (iii) an efficient architecture for real-time interaction. We show an overview of training in Fig. 3 and 4.
@@ 3.1 Problem Formulation
Let $t$ be a segmentation task consisting of image and segmentation pairs, $\{(x^{t},y^{t})_{j}\}_{j=1}^{N}$. At step $i$, given an image $x^{t}$, a set of user interactions $u_{i}$, and previous prediction $\hat{y}^{t}_{i-1}$, we learn function $f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$ with parameters $\theta$ that produces a segmentation $\hat{y}_{i}$. The set of interactions $u_{i}$, which may include positive or negative scribbles, positive or negative clicks, and bounding boxes, is provided by a user who has access to the image $x^{t}$ and previous prediction $\hat{y}_{i-1}^{t}$.
We minimize the difference between the true segmentation $y^{t}$ and each of the iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$,
$\mathcal{L}(\theta;\mathcal{T})=\mathbb{E}_{t\in\mathcal{T}}\left[\mathbb{E}_{(x^{t},y^{t})\in t}\left[\sum_{i=1}^{k}\mathcal{L}_{Seg}\left(y^{t},f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})\right)\right]\right],$ (1)
@@
During training, we sample a task $t$ from training task collection $\mathcal{T}$, from which we sample an image $x^{t}$, and segmentation map $y^{t}$. We simulate a possible set of interactions $u_{i}$ based on $y^{t}$, and predict $\hat{y}_{i}^{t}=f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$. We simulate the next set of interactions $u_{i+1}$ based on the error region $\varepsilon_{i}^{t}=y^{t}-\hat{y}_{i}^{t}$, and repeat for $k$ iterations. In the following sections, we describe the core aspects of the framework: strategies for simulating $u_{i}$, collecting $\mathcal{T}$, sampling $(x^{t},y^{t})$, building $f_{\theta}$, and optimizing $\mathcal{L}(\theta;\mathcal{T})$.
@@ 3.2 Prompt Simulation
To enable a practical, easy-to-use model, we encourage robustness to different types of user interactions $u_{i}$. We introduce algorithms for simulating scribbles, clicks, and bounding box inputs during training. Each scribble and click strategy (Fig. 2) can be applied to a ground truth segmentation label $y^{t}$ to simulate positive interactions, or to the background $1-y^{t}$ to simulate negative interactions. We simulate positive and negative correction scribbles or clicks by applying the same strategies to the false negative region $\varepsilon_{i}^{t}>0$ and false positive region $\varepsilon_{i}^{t}<0$ for error region $\varepsilon_{i}^{t}$.
Scribbles. Given binary mask $y\in\{0,1\}^{h\times w}$, we simulate a scribble mask $s\in[0,1]^{h\times w}$ by first generating clean scribbles, and then corrupting them to account for user behavior and variability. We illustrate this process in Fig. 2 and Sec. B.1.1. To generate the clean scribbles, we start with one of these strategies:
@@
    Line Scribbles: We draw random lines by connecting two end points sampled from $\{(u,v)|y_{uv}=1\}$ (Fig. 8).
@@
    Centerline Scribbles: We simulate scribbles in the center of label $y$ using a thinning algorithm [147] that reduces the label to a 1-pixel wide skeleton (Fig. 9).
@@
    Contour Scribbles: We simulate a rough contour of the desired segmentation within the boundaries of the mask (Fig. 10). We first blur the mask to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{y}<h$ sampled in some intensity range $h\sim U[\tilde{y}_{min},\tilde{y}_{max}]$ and extract a contour inside the boundary of the mask.
Figure 2: Simulated scribbles and clicks. Positive interactions (green) are simulated on the segmentation label $y^{t}$ (blue), while negative interactions (red) are simulated on the background $1-y^{t}$. Scribble thickness is enlarged for visual clarity.
To limit the size and complexity of the centerline and contour scribbles, we apply a random mask that breaks the scribbles into smaller parts. We generate the random mask by sampling a smooth noise image $p$, where each pixel is sampled independently from $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$, applying Gaussian blur, and then thresholding it at $\mu_{p}$. We warp the resulting scribble mask $s$ using a random deformation field $\phi$ to vary the scribble shape and thickness. We ensure the resulting scribble is consistent with mask $y$, by multiplying the warped scribble mask $s\circ\phi$ by $y$.
Clicks. Given a binary mask $y\in\{0,1\}^{h\times w}$, we simulate $n\sim U[n_{min},n_{max}]$ clicks at a time using one of three strategies, illustrated in Fig. 2:
@@
    Center clicks: We sample clicks from the set of points at the center of each disconnected component of the label. First, we create a multi-label mask $m\in\{1,\dots,C\}^{h\times w}$ identifying the $C$ components of label $y$. We identify the center of each components using the euclidean distance transform [123, 140].
@@
    Interior border region clicks: We sample clicks from a border region inside the boundary of the mask (Fig. 11). We first blur the mask $y$ to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{y}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{y}_{min},\tilde{y}_{max})$ are thresholds sampled in some intensity range.
Since users are inclined to spread out their clicks, we impose a minimum separation of a few pixels between random clicks and border region clicks.
Bounding Boxes. We compute the minimum bounding box that encloses the label $y^{t}$, and enlarge each dimension by $r\sim U[0,20]$ pixels to account for human variability.
Iterative Training (Fig. 3). During the first step ($i=1$), we sample the combination of interactions and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The initial interactions $u_{1}$ are simulated using the ground truth label $y^{t}$. In subsequent steps, correction scribbles or clicks are sampled from the error region $\varepsilon_{i-1}^{t}$ between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth $y^{t}$. Since a user can make multiple corrections in each step, we sample $n_{cor}\sim U[n_{min},n_{max}]$ corrections (scribbles or clicks) per step.
Figure 3: Training. We simulate $k$ consecutive steps of interactive segmentation. Given an image segmentation pair $(x^{t},y^{t})$, we first simulate a set of initial interactions $u_{1}$, which may contain bounding boxes, clicks, and/or scribbles. We predict segmentation $\hat{y}_{1}^{t}:=f_{\theta}(x^{t},u_{1},\hat{y}_{0}^{t})$ where the initial prediction $\hat{y}^{t}_{0}$ is set to zeros. In the second step, we simulate corrections using the error region $\varepsilon_{1}$ between the previous prediction $\hat{y}_{1}$ and ground truth $y$, and add them to the set of initial interactions $u_{1}$ to get $u_{2}$. We predict segmentation $\hat{y}_{2}:=f_{\theta}(x,u_{2},\hat{y}_{1})$ and repeat to produce a series of $k$ iterative predictions, $\hat{y}_{1},\dots,\hat{y}_{k}$. We learn $\theta$ to minimize $\sum_{i=1}^{k}\mathcal{L}_{seg}(y^{t},\hat{y}_{i})$, the sum of losses between the target segmentation $y^{t}$ and iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$.